OpenStudy (anonymous):

@ if A^-1(inverted) matrix exists and B matrix and S matrix prove these AB=BA and S^2 =B prove that (A^-1SA)^2=B

OpenStudy (jamesj):

1. A is invertible 2. AB = BA 3. S^2 = B Show that \( (A^{-1}SA)^2 = B \). So just start calculating: \[ (A^{-1}SA)^2 = A^{-1}SAA^{-1}SA = ... \]

OpenStudy (zarkon):

nice translation of the question ... :)

OpenStudy (jamesj):

Thanks. ;-) I guess I shouldn't be surprised, but it is interesting how often clarifying the question to "well posed" is necessary on here.

OpenStudy (anonymous):

so in the end it is ...=I*B am I right ?

OpenStudy (jamesj):

and IB = B, by definition of I